A room heater is rated at 12 amps and 120vac. Electricity cost $0.12 per kWh. How much would it cost to run this for one month if it stayed on all of the time
1
Expert's answer
2020-04-23T12:43:49-0400

The power of the heater is "P = I U = 12 A \cdot 120 V = 1440W = 1.44kW".

The duration of a 30 day month is "t = 30 \cdot 24 h = 720 h", hence the amount of energy the heater will consume if stayed on all the time is "1.44 kW \cdot 720 h = 1036.8 kW \cdot h".

This will cost "0.12 \$ \cdot 1036.8 kW \cdot h \approx 124.42 \$."